OpenStudy (anonymous):

how would i do this y=2x +3 containing (-2,2) finding the equation of the line that contain these given point and is parallel to the given line graph the lines.

OpenStudy (tyteen4a03):

Substitute (-2, 2) into the equation. If both sides are equal then that point lies on the line.

OpenStudy (anonymous):

so 2=-6x+3?

OpenStudy (anonymous):

You are going to use Point/Slope line format -- It looks like this: y-y1 = m(x-x1), where m is the slope and (x1, y1) is the point on the line. Your slope is going to be perpendicular to the given line. That means the slope will be the negative inverse of the slope of y=2x+3. The m of this line is 2; its negative inverse is -1/2. So, you want the equation of a line with a slope of -1/2, that goes through the point (-2,2). Plus (-1/2) in for slope, 2 in for y1, and -2 in for x1: y-2=-1/2(x+2)

OpenStudy (anonymous):

OOPs - you want the parallel slope, not perpendicular -- so the slope would just be 2: y-2=2(x+2)

OpenStudy (anonymous):

so I was suppose to use y-y1=something(x-x1) form and then plug them into equation?

OpenStudy (anonymous):

yes - where "something" is the slope. Parallel lines have the same slope, so you get that from the equation you're given.

OpenStudy (anonymous):

so y=2x?

OpenStudy (anonymous):

I get y-2=2(x+2) y-2 = 2x + 4 y = 2x+6 y = 2x + 6

OpenStudy (anonymous):

oh i thought 2 canceled each other out because when i tok it from 1 side forgot to turn it to positive

OpenStudy (anonymous):

Easy mistake to make :)

OpenStudy (anonymous):

well thank you ima try apply this to other problems too.

OpenStudy (anonymous):

Good luck and happy mathing!